OpenStudy (buttermequeasy):

Solve: −2(x − 5) = −4

OpenStudy (anonymous):

-2x +10 = -4 -2x = -14 x = 7

OpenStudy (anonymous):

x= 7

OpenStudy (anonymous):

use distributive propety